Solve using Gaussian Elimination:
3x + 4y = 7
-5x + 2y = 10

Answer by JimboP1977(308)

3x + 4y = 7 (1)
-5x + 2y = 10 (2)
(2) *2
Gives -10x + 4y = 20 (3)
(3) - (1)
-13x = 13
x = -1
y = 2.5